OU 18. (1-x) + x - y = 0, 7(0) = -3, y (0) = 2 s eL TUDOM 9. (a) By making the change of variable x-1 = 1 and assuming that problem 19
Answer #1


dt Page-1 n=0 *-=t - Edt doo ; (07-1)=(26-1)(3+1) dy = dy.dt dy =t(t+2) dz de la doce de leurs et des equation becomes Y+ty3.2 C3 +260 =0 4.3 C4 +1.C + Co +2C,=0 12 C4 +36 +60=0 n(n-1) Cnt (n-3) Cn-3 + Cn-4+2 Cn-50 min-1) Cn +(2-1) G2-3 + Cn-450 10